OpenStudy (anonymous):

log(10^6(x^2)) I am really bad with logarithms..

OpenStudy (unklerhaukus):

\[\log(10^6(x^2))\]

OpenStudy (unklerhaukus):

are you trying to simplify?

OpenStudy (unklerhaukus):

These will help\[\boxed{\log_bxy=\log_b x+\log_by}\]\[\boxed{\log_b x^n=n\log_bx}\]\[\boxed{\log a=\log_{10}a}\]\[\boxed{\log_b b=1}\]